OpenStudy (anonymous):

Can someone please help me with this question? Firm 1 has a capital structure with 20% debt and 80% equity. Firm 2’s capital structure consists of 50% debt and 50% equity. Both firms pay 7% annual interest on their debt. Finally suppose both firms have invested in assets worth $100 million. Calculate the return on equity (ROE) for each firm assuming the following: a. The return on assets is 3% b. The return on assets is 7% c. The return on assets is 11% What general pattern do you observe?

OpenStudy (anonymous):

If i may try to answer the question, (cmiiw): first thing first, the ROA = 3% means, the Return compared to Total Assets = 3:100 which gives us 3 million as a Return. Usually there are 2 approaches to compute ROA, one is using the net income, and the other is using the EBIT (or so i was told) condition 1 (ROA = Net Income) Net Income = 3,000,000 while Assets = 100,000,000. We know that Total Assets are equal to Total Liabilities + Total Equities. Given the proportion of the Debt&Equity, we can compute the ROE for each Firm: Firm 1 with 20% Debt -80% Equity, that means: Total Debt = 20 million Total Equity = 80 million With the Net Income of 3,000,000 the ROE will be 3 million / 80 million = 3.75% Firm 2 with 50% Debt -50% Equity, that means: Total Debt = 50 million Total Equity = 50 million With the Net Income of 3,000,000 the ROE will be 3 million / 50 million = 6% however, given the information about the interest rate of 7%, i assume that the ROA is computed using the EBIT, hence the computation will be slight different, which is: 3,000,000 - Interest Expense - Tax (assuming there is no tax) = Net Income so for the Firm 1 3,000,000 - (7% x 20,000,000) = 2,600,000 with that amount of NI, the ROE is now: 2,600,000 / 80,000,000 = 3.25% for the Firm 2 3,000,000 - (7% x 50,000,000) = -500,000 and with that amount of NI, the ROE is: -500,000 / 50,000,000 = (10%) and i believe the other conditions will use the same approach... about the intepretation, i honestly can't do the intepretation since i'm not quite sure which assumption of mine, is correctly used in the calculation.. hence i'll leave the rest to you ^^ i hope i'm not misleading you..